### v1

Abstract
Quadrupedal locomotion using deep reinforcement learning (DRL) is an enabling technology for terrain negotiation, stable mobility, and real-world deployment, and is a key capability in legged robotics. Existing approaches either train blind agents with heavy domain randomization, which sacrifices foresight, or adopt cross-modal fusion architectures that have their own limitations, where recurrent neural networks struggle to model long-term dependencies due to vanishing gradients, hierarchical designs complicate optimization and propagate errors across levels, and Transformer-based fusion incurs quadratic memory and computational cost from self-attention. To address these challenges, we introduce LocoMamba, a vision-driven cross-modal DRL framework built on selective state-space models, specifically leveraging Mamba, that achieves near-linear-time sequence modeling, effectively captures long-range dependencies, and enables efficient training with longer sequences. First, we embed proprioceptive states with a multilayer perceptron and patchify depth images with a lightweight convolutional neural network, producing compact tokens that improve state representation. Second, stacked Mamba layers fuse these tokens via near-linear-time selective scanning, reducing latency and memory footprint, remaining robust to token length and image resolution, and providing an inductive bias that mitigates overfitting. Third, we train the policy end-to-end with Proximal Policy Optimization under terrain and appearance randomization and an obstacle-density curriculum, using a compact state-centric reward that balances progress, smoothness, and safety. We evaluate our method in challenging simulated environments with static and moving obstacles as well as uneven terrain. Compared with state-of-the-art baselines, our method achieves higher returns and success rates with fewer collisions, exhibits stronger generalization to unseen terrains and obstacle densities, and improves training efficiency by converging in fewer updates under the same compute budget.
keywords:
[…]
{highlights}
State-of-the-art LocoMamba: To the best of our knowledge, this is the first vision-driven cross-modal DRL framework for quadrupedal locomotion that utilizes the selective state-space model Mamba as the fusion backbone, enabling foresightful and efficient control.
Effective proprioception and depth encoders: We introduce a compact input representation in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and look-ahead while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
Efficient cross-modal Mamba fusion backbone: We fuse encoded tokens from proprioception and depth using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory growth. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
Robust end-to-end RL training scheme: We train the policy with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances progress, smoothness, efficiency, and safety, enabling stable learning and consistent performance.
Comprehensive evaluation: We conduct extensive experiments with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, success rate, and collision rate, along with faster convergence under the same compute budget.
1 Introduction
Quadrupedal robots provide mobility in environments where wheeled platforms are ineffective, such as stairs, rubble, soft or deformable substrates, and cluttered indoor or outdoor settings, enabling applications in inspection, disaster response, agriculture, and planetary exploration (Fan et al., 2024). Robust locomotion control is therefore a foundational capability for practical quadrupedal systems, underpinning safe navigation, dependable mission execution, and reliable operation across diverse terrains and disturbance conditions (Carpentier and Wieber, 2021).
Current approaches to quadrupedal locomotion control fall into two broad categories. Classical model-based pipelines depend on accurate dynamics and terrain models, explicit contact planning, and extensive parameter tuning, which constrain scalability and complicate deployment (Xin et al., 2021). Learning-based methods mitigate this burden and enable end-to-end policies that couple perception with control (Ha et al., 2025). Within this paradigm, deep reinforcement learning has emerged as the dominant approach because it optimizes task-driven closed-loop policies through interaction with the environment and produces robust, adaptive behaviors (Zhang et al., 2022).
DRL methods have substantially advanced quadrupedal locomotion, enabling traversal of uneven terrain (Xie et al., 2021) and operation under challenging conditions such as mud, snow, and running water (Lee et al., 2020). However, most studies continue to train blind controllers that rely only on proprioception and achieve stability primarily through large-scale simulation and domain randomization. Despite their robustness, blind agents lack foresight because they receive no exteroceptive input, so they react only after contact and struggle to avoid obstacles proactively or to plan foot placement on irregular ground (Xie et al., 2022).
On the other hand, cross-modal fusion with visual input has gained traction. Vision complements proprioception by providing look-ahead, enabling early perception of distant obstacles, anticipation of terrain changes before contact, and timely trajectory adjustments that reduce collisions and improve foot placement (Han et al., 2025). Nevertheless, current fusion architectures exhibit notable trade-offs. Recurrent models often suffer from vanishing gradients and limited capacity for long-horizon dependencies (Li et al., 2024). Hierarchical designs complicate optimization and propagate errors across levels (Jain et al., 2019). Transformer-based fusion, although expressive, incurs quadratic memory use and computational cost with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency (Yang et al., 2021).
In this paper, we propose LocoMamba, a vision-driven end-to-end DRL framework that uses Mamba as a selective state-space model (SSM) backbone to enable efficient quadrupedal locomotion. A lightweight multilayer perception (MLP) embeds proprioceptive states to provide accurate estimates of the robot state for immediate reaction, while a compact convolutional neural network (CNN) patchifies depth images to supply look-ahead for negotiating uneven terrain and large obstacles. Stacked Mamba layers fuse tokens through selective state-space scanning with near-linear-time complexity, replacing quadratic self-attention with recurrent updates that lower latency and memory use. The streaming formulation accommodates variable-length inputs, enabling longer visual context and higher token resolution, while input-gated, exponentially decaying dynamics provide a regularizing inductive bias that mitigates overfitting. Policies are trained with Proximal Policy Optimization (PPO) under terrain and appearance randomization together with an obstacle-density curriculum, which broadens environmental coverage and progressively increases difficulty to stabilize on-policy learning. The reward is compact and state-centric, encouraging progress and accurate tracking, promoting smooth and efficient actuation through torque and acceleration regularization, and enforcing safety through collision, fall, and foot-clearance terms.
The main contributions of this paper are summarized as follows:
  1. [1)]
  2. 1.
    State-of-the-art LocoMamba: To the best of our knowledge, this is the first vision-driven cross-modal DRL framework for quadrupedal locomotion that utilizes the selective state-space model Mamba as the fusion backbone, enabling foresightful and efficient control.
  3. 2.
    Effective proprioception and depth encoders: We introduce a compact input representation in which an MLP embeds proprioceptive states and a CNN patchifies depth images into spatial tokens tailored for Mamba-based fusion. This design provides immediate state estimates and look-ahead while reducing sensitivity to appearance variation, thereby improving computational efficiency and training stability.
  4. 3.
    Efficient cross-modal Mamba fusion backbone: We fuse encoded tokens from proprioception and depth using stacked Mamba layers via selective state-space scanning, achieving near-linear time and memory growth. The backbone supports long-horizon modeling, remains robust to token length and image resolution, and provides a regularizing inductive bias through input-gated, exponentially decaying dynamics.
  5. 4.
    Robust end-to-end RL training scheme: We train the policy with PPO using Mamba-fused cross-modal features, complemented by terrain and appearance randomization and an obstacle-density curriculum. A compact, state-centric reward balances progress, smoothness, efficiency, and safety, enabling stable learning and consistent performance.
  6. 5.
    Comprehensive evaluation: We conduct extensive experiments with static and moving obstacles and uneven terrain, and demonstrate consistent gains over state-of-the-art (SOTA) in terms of return, success rate, and collision rate, along with faster convergence under the same compute budget.
The remainder of this paper is organized as follows. Section 2 reviews related work on blind and cross-modal quadruped locomotion. Section 3 presents the LocoMamba methodology, including the problem formulation, proprioception and depth encoders, and the Mamba fusion backbone with training objectives. Section 4 describes the experimental setup and evaluation protocol. Section 5 presents the experimental evaluation, covering the main results and ablations, as well as extended validation of efficiency and generalization. Finally, Section 6 concludes the paper and outlines directions for future research.
2 Related Work
2.1 Learning-Based Quadrupedal Locomotion
Research on quadrupedal locomotion has evolved along two principal paradigms: rule-based control and learning-based approaches. Early rule-based controllers relied on template dynamics and heuristic regulation of balance, leg compliance, and foot placement, establishing fundamental principles for dynamic locomotion (Miura and Shimoyama, 1984; Liu and Ding, 2014; Habu et al., 2018; Bledt et al., 2018). Subsequent rule-based systems formalized locomotion as constrained optimal control, including trajectory optimization over centroidal or full-body dynamics and model predictive control with explicit contact constraints and task hierarchies (Grandia et al., 2019; Amatucci et al., 2024; Elobaid et al., 2025). For example, (Di Carlo et al., 2018) developed real-time MPC and task-space whole-body control on torque-controlled quadrupeds, demonstrating stable trotting, stair traversal, and robust disturbance rejection. (Ding et al., 2019) refined MPC cost design and contact modeling to improve tracking and compliance under high-frequency control budgets. (Carius et al., 2019) studied trajectory optimization and contact-aware planning for foothold selection and force allocation, improving reliability on structured terrain. However, these approaches require accurate models of the environment and system dynamics and substantial manual tuning, thereby limiting their applicability in complex and variable settings.
On the other hand, model-free RL can learn general policies for challenging conditions (Li et al., 2021; Margolis et al., 2024; Bussola et al., 2025; Ha et al., 2025). For example, (Hwangbo et al., 2019) trained torque-level policies with dynamics randomization and privileged learning to achieve fast, robust trotting on uneven terrain. (Tan et al., 2018) used large-scale domain randomization and dynamics perturbations to transfer simulation-trained policies to real quadrupeds with strong disturbance rejection. (Kumar et al., 2021) introduced rapid motor adaptation that augments a proprioceptive policy with an online adaptation module, enabling quick recovery under payload shifts and terrain changes. However, most approaches are proprioception-only and thus blind to exteroceptive cues, which limits foresight for obstacle negotiation and precise foothold planning. In this work, we fuse visual and proprioceptive inputs to obtain a richer state representation, enabling the policy to anticipate terrain changes and plan trajectories online while maintaining continuous motion.
2.2 Vision-Driven RL for Quadrupedal Locomotion
To extend RL beyond state-only inputs, a growing body of work leveraged visual observations for locomotion control (Han et al., 2025). For example, Yu et al. (2021) trained a vision-based controller that incorporated exteroceptive input directly into the RL loop and achieved traversal of uneven terrain and complex obstacles. Duan et al. (2024) learned a controller in simulation with heightmap observations and then trained a depth-to-heightmap predictor from depth and state histories, enabling vision-guided locomotion with transfer to hardware on challenging terrains. Complementing these efforts, Fahmi et al. (2022) proposed the ViTAL framework, which used vision for terrain-aware planning by decomposing decision making into foothold selection and pose adaptation, and improved safety and reliability on irregular surfaces. Hierarchical formulations were also explored; for example, Jain et al. (2019) employed hierarchical RL in which high-level visual policies guided low-level motor control with sensor inputs.
Despite these advances, common fusion strategies involved trade-offs. Simple concatenation underused spatial structure, and hierarchical decompositions increased optimization complexity and risked error propagation across levels (Singh et al., 2022). In this work, we avoid complex hierarchical designs and adopt a simple end-to-end schema that fuses proprioceptive and depth inputs within a single policy trained with PPO, enabling visual look-ahead while maintaining a streamlined learning pipeline.
2.3 State-Space Models and Cross-Modal RL
Cross-modal RL has attracted significant attention. Typical fusion backbones include recurrent neural networks (RNNs) and Transformers. For example, (Xiao et al., 2024) designed an egocentric visuomotor RL framework that fused egocentric depth and proprioception with long short-term memory (LSTM) units, enabling a quadruped to negotiate obstacles and traverse cluttered scenes in the real world. (Lai et al., 2024) proposed a gated recurrent unit (GRU)–based world-model formulation that fuses depth images and proprioceptive readings into a recurrent latent state for visual legged locomotion, improving control under partial observability. (Yang et al., 2021) introduced a cross-modal Transformer policy that ingests depth tokens and proprioceptive features end-to-end, achieving stronger terrain anticipation and sim-to-real transfer than blind baselines; subsequent work scaled such tokenized proprioception-vision Transformer policies to heterogeneous robot datasets to improve generalization. However, RNNs often suffer from vanishing gradients and limited long-horizon capacity, which complicates optimization in extended sequences. Transformer models offer strong expressivity but incur quadratic memory and computation with the number of tokens, which constrains sequence length, spatial resolution, and training efficiency.
SSMs offer an alternative by updating a compact recurrent state per token and scaling near-linearly with sequence length. Building on the S4 model (Gu et al., 2021), which demonstrated efficient long-range dependency modeling on long-sequence benchmarks, (Gu and Dao, 2023) introduced Mamba , a selective SSM whose input-dependent parameters and hardware-aware parallel scan yield linear-time sequence modeling with favorable throughput and memory characteristics across modalities.
Mamba-style SSMs have also been explored for multimodal fusion. AlignMamba augments a multimodal Mamba backbone with token-level optimal-transport alignment and a global distributional alignment objective to improve cross-modal consistency (Li et al., 2025). COMO (Cross-Mamba Interaction with Offset-Guided Fusion) addresses sensor misalignment for multimodal object detection while retaining the efficiency benefits of selective scanning (Liu et al., 2026). AV-Mamba applies selective SSMs to audio-visual question answering and reports improvements over Transformer baselines (Huang et al., 2024). FusionMamba (Xie et al., 2024) and DepMamba (Ye et al., 2025) further demonstrate SSM-based gains in multimodal image fusion and audio-visual affect analysis. To the best of our knowledge, this is the first work to use cross-modal Mamba for quadrupedal locomotion.
3 Methodology
Figure 1 presents the overall architecture of LocoMamba. The pipeline comprises three components trained end to end. First, proprioceptive states and depth images are encoded into a compact tokenized latent space: an MLP maps the proprioceptive vector to a state token, and a lightweight CNN patchifies the depth image into spatial tokens. Second, the concatenated token stream is fused by stacked Mamba SSM layers via selective state-space scanning, which updates a compact recurrent state per token, yields near linear time and memory scaling, accommodates variable token counts and resolutions, and preserves long-horizon temporal context. Third, policy and value heads consume the fused features and are optimized with PPO under a compact state-centric reward that balances progress, smoothness and efficiency, and safety.
Figure 1: Overall architecture of LocoMamba. Proprioception and depth are encoded into tokens by MLP/CNN, fused by a Mamba SSM backbone, and optimized with PPO.
3.1 Proprioception–Depth Input Encoding
As described in Section 1, we consider an agent that uses both proprioceptive state and depth image for decision making. At time $t$, the observation is
$o_{t}=\{\,s^{\text{prop}}_{t},\;I^{\text{depth}}_{t}\,\},$ (1)
where $s^{\text{prop}}_{t}\in\mathbb{R}^{D_{p}}$ denotes the proprioceptive vector and $I^{\text{depth}}_{t}\in\mathbb{R}^{H\times W}$ denotes a first-person depth image that captures obstacles and terrain ahead of the robot. Temporal aggregation of the proprioceptive vector and the four most recent depth frames is handled by the fusion backbone described in Section 3.2.
Encoders and tokenization.
We encode each modality with a lightweight, domain-specific network and unify them in a shared latent space. The proprioceptive vector is mapped by an MLP to a compact token:
$z^{\text{prop}}_{t}=f_{\text{MLP}}(s^{\text{prop}}_{t}),\quad z^{\text{prop}}_{t}\in\mathbb{R}^{d_{p}}.$ (2)
The depth image is processed by a compact CNN to produce a spatial feature map:
$Z^{\text{vis}}_{t}=f_{\text{CNN}}(I^{\text{depth}}_{t}),\quad Z^{\text{vis}}_{t}\in\mathbb{R}^{N\times d_{v}}.$ (3)
With patch size $P$, the number of spatial tokens is
$N=\frac{H}{P}\cdot\frac{W}{P}.$ (4)
To form a single cross-modal stream, we project both embeddings to a common width $d$. The projected proprioceptive token is
$\tilde{z}^{\text{prop}}_{t}=W_{p}\,z^{\text{prop}}_{t},\quad\tilde{z}^{\text{prop}}_{t}\in\mathbb{R}^{d}.$ (5)
The projected visual tokens are
$\tilde{Z}^{\text{vis}}_{t}=W_{v}\,Z^{\text{vis}}_{t},\quad\tilde{Z}^{\text{vis}}_{t}\in\mathbb{R}^{N\times d}.$ (6)
where $W_{p}$ and $W_{v}$ are learned projections and $N=N_{h}N_{w}$ denotes the number of spatial tokens. We then concatenate the modalities into one token sequence
$U_{t}=\big{[}\tilde{z}^{\text{prop}}_{t};\tilde{Z}^{\text{vis}}_{t}\big{]},\quad U_{t}\in\mathbb{R}^{(1+N)\times d}.$ (7)
Finally, we add learned position indices and modality tags, followed by per-token layer normalization, to obtain
$\hat{U}_{t}=\mathrm{LN}\!\left(U_{t}+E_{\text{pos}}^{\text{spat}}+E_{\text{mod}}\right).$ (8)
where $E_{\text{pos}}^{\text{spat}}\in\mathbb{R}^{(1+N)\times d}$ encodes spatial indices (the proprioceptive token receives a distinct index), and $E_{\text{mod}}\in\mathbb{R}^{(1+N)\times d}$ distinguishes proprioceptive and visual tokens.
3.2 State Space Modelling with Mamba
We fuse the cross-modal token sequence $\hat{U}_{t}$ with a stack of Mamba SSM layers (Fig. 2). Each layer scans the token stream with input-dependent state updates and carries a compact recurrent state across tokens and time steps, enabling efficient long-horizon modeling.
Figure 2: Schematic of a Mamba SSM block.
Per-layer selective SSM.
Let $u_{t,k}\in\mathbb{R}^{d}$ denote the $k$-th token at time $t$ from $\hat{U}_{t}$ (k=1k{=}1 is the proprioceptive token, $k{=}2\ldots 1{+}N$ are visual tokens in raster order). A Mamba layer maintains a hidden state $x_{t,k}\in\mathbb{R}^{h}$ and computes
$x_{t,k+1}\;=\;\bar{A}_{t,k}(u_{t,k})\,x_{t,k}\;+\;\bar{B}_{t,k}(u_{t,k})\,u_{t,k},$ (9)
$y_{t,k}\;=\;\bar{C}_{t,k}(u_{t,k})\,x_{t,k}\;+\;\bar{D}_{t,k}(u_{t,k})\,u_{t,k},$ (10)
where $\bar{A}_{t,k},\bar{B}_{t,k},\bar{C}_{t,k},\bar{D}_{t,k}$ are input-gated, token-dependent parameters produced by lightweight affine transforms of $u_{t,k}$. The scan proceeds over $k{=}1\ldots(1{+}N)$, and the initial state at each time step is carried from the previous step, $x_{t,1}\leftarrow x_{t-1,\,1+N}$. The output tokens of the layer are $Y_{t}=\{y_{t,k}\}_{k=1}^{1+N}\in\mathbb{R}^{(1+N)\times d}$. We set $h$ and $d$ equal via a learned projection when needed.
Stacked fusion backbone and complexity.
We stack $L_{\text{m}}$ Mamba layers with residual connections and layer normalization:
$H^{(\ell+1)}_{t}\;=\;\mathrm{LN}\!\left(Y^{(\ell)}_{t}+H^{(\ell)}_{t}\right),\qquad\ell=0,\ldots,L_{\text{m}}-1,$ (11)
with $H^{(0)}_{t}\!=\!\hat{U}_{t}$ and $Y^{(\ell)}_{t}$ computed by (9)–(10) using inputs $H^{(\ell)}_{t}$. The selective SSM update yields near-linear time and memory in the token count (1+N)(1{+}N), in contrast to the quadratic cost of global self-attention, and naturally supports streaming over time by reusing the carried state $x_{t,1}$.
Projection head for policy inputs.
To obtain a compact feature for control, we perform modality-aware pooling on the final layer outputs $H^{(L_{\text{m}})}_{t}$. Let $y^{\text{prop}}_{t}$ be the proprioceptive token and $\{y^{\text{vis}}_{t,i}\}_{i=1}^{N}$ the visual tokens from $H^{(L_{\text{m}})}_{t}$. We compute
$\bar{y}^{\text{vis}}_{t}=\frac{1}{N}\sum_{i=1}^{N}y^{\text{vis}}_{t,i},\qquad h_{t}=f_{\text{head}}\!\left(\big{[}\,y^{\text{prop}}_{t}\,;\,\bar{y}^{\text{vis}}_{t}\,\big{]}\right)\in\mathbb{R}^{d_{h}},$ (12)
where $f_{\text{head}}$ is a small MLP. The fused feature $h_{t}$ parameterizes the policy and value functions in Section 3.3.
Positional coding and robustness.
Spatial indices are encoded in $E_{\text{pos}}^{\text{spat}}$ and preserved throughout the scan. When temporal indices are used, an additional code $E_{\text{pos}}^{\text{time}}$ is added to the tokens before the first Mamba layer. The recurrent state $x_{t,k}$ provides a causal, exponentially decaying memory that promotes temporally consistent features and improves robustness to variable token counts and resolutions, while keeping latency and memory use low.
3.3 Policy Optimization with PPO
Markov Decision Process (MDP).
We model locomotion as a discounted MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$ with discount $\gamma\in(0,1)$. At time $t$, the agent receives observation $o_{t}$ and obtains a fused feature $h_{t}=f_{\text{fuse}}(o_{t})\in\mathbb{R}^{d_{h}}$ from the Mamba backbone (Sections 3.2). The objective is to maximize the expected discounted return
$J(\theta)=\mathbb{E}\!\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t}\right],$ (13)
where the expectation is over on-policy trajectories, $T$ is the horizon, and $r_{t}=r(s_{t},a_{t})$.
The action is also a 12-dimensional vector that controls the change of all the joint angles. We use a Gaussian policy in the unconstrained space with squashing to actuator limits:
$\tilde{a}_{t}\sim\mathcal{N}\!\big{(}\mu_{\theta}(h_{t}),\,\operatorname{diag}(\sigma_{\theta}^{2}(h_{t}))\big{)},\qquad a_{t}=a_{\max}\tanh(\tilde{a}_{t}).$ (14)
where $\mu_{\theta}(\cdot)$ and $\sigma_{\theta}(\cdot)$ are outputs of the policy head.
Policy and value parameterization.
Two MLP heads consume $h_{t}$: the first produces $\mu_{\theta}(h_{t})$ and $\log\sigma_{\theta}(h_{t})$, which parameterize the policy $\pi_{\theta}(a_{t}\mid h_{t})$, and the second outputs $V_{\phi}(h_{t})$ as the state-value estimate.
PPO objective and estimation.
We optimize the policy with PPO using the clipped surrogate. With likelihood ratio
$\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid h_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid h_{t})},$ (15)
the policy loss is
$\mathcal{L}_{\text{clip}}(\theta)=\mathbb{E}\Big{[}\min\big{(}\rho_{t}(\theta)A_{t},\ \operatorname{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)\,A_{t}\big{)}\Big{]},$ (16)
where $\epsilon>0$ is the clip parameter and $A_{t}$ are advantages. Advantages use generalized advantage estimation (GAE),
$\delta_{t}=r_{t}+\gamma V_{\phi}(h_{t+1})-V_{\phi}(h_{t}),\qquad A_{t}=\sum_{l=0}^{T-1-t}(\gamma\lambda)^{l}\,\delta_{t+l},$ (17)
where $\lambda\in[0,1]$ controls bias–variance trade-offs. The critic and entropy terms are
$\mathcal{L}_{V}(\phi)=\mathbb{E}\big{[}(V_{\phi}(h_{t})-\hat{R}_{t})^{2}\big{]},\qquad\mathcal{H}_{t}=\mathcal{H}\!\big{[}\pi_{\theta}(\cdot\mid h_{t})\big{]},$ (18)
where $\hat{R}_{t}$ is the empirical return and $\mathcal{H}$ denotes entropy. The total loss is
$\mathcal{J}(\theta,\phi)=-\,\mathcal{L}_{\text{clip}}(\theta)+\beta_{V}\,\mathcal{L}_{V}(\phi)-\beta_{H}\,\mathbb{E}[\mathcal{H}_{t}],$ (19)
where $\beta_{V},\beta_{H}\geq 0$ weight value and entropy terms.
4 Implementation
4.1 Computing Environment
We conduct all experiments on a cloud workstation equipped with an Intel Xeon Gold 6430 CPU (128 cores, 2.1 GHz base clock) and an NVIDIA GeForce RTX 4090 GPU (24 GB). The operating system is Ubuntu 22.04. We use PyBullet for physics simulation and implement all models in Python with PyTorch 2.4.1.
4.2 Simulation Environment
We evaluate across six simulated environments (Yang et al., 2021) that vary terrain difficulty, obstacle geometry, and the presence of collectible spheres (used as sparse task bonuses):
  • Wide Obstacle (Wide Obs.): flat terrain with wide cuboid obstacles; no spheres.
  • Thin Obstacle (Thin Obs.): flat terrain with numerous thin cuboid obstacles; no spheres.
  • Wide Obstacle & Spheres (Wide Obs. & Sph.): as Wide Obs., with collectible spheres for bonus reward.
  • Thin Obstacle & Spheres (Thin Obs. & Sph.): thin obstacles plus collectible spheres on flat terrain.
  • Moving Obstacle (Moving Obs.): similar to Thin Obs., but obstacles move in random directions with low-frequency updates.
  • Mountain: rugged terrain with a goal at the summit (progress toward the goal).
Figure 3 shows representative examples. Unless noted, obstacles and spheres are randomized at episode reset and remain static; only Moving Obs. updates obstacle positions during an episode.
Figure 3: Simulated environments. Panels (a) to (f): (a) Wide Obs.; (b) Thin Obs.; (c) Wide Obs. & Sph.; (d) Thin Obs. & Sph.; (e) Moving Obs.; (f) Mountain. Obstacles and spheres are randomized at reset and remain static except in (e).
4.3 RL MDP Details
Observation space.
The observation at time $t$ comprises a 93-dimensional proprioceptive vector $s^{\text{prop}}_{t}$ (IMU readings, local joint rotations, and the actions issued over the previous three time steps) and a stack of the four most recent dense depth images $\big{[}I^{\text{depth}}_{t-3},\,I^{\text{depth}}_{t-2},\,I^{\text{depth}}_{t-1},\,I^{\text{depth}}_{t}\big{]}$, each with resolution $64\times 64$.
Action space.
The action is a 12-dimensional vector that controls the change of all the joint angles.
Reward function.
We use a unified reward that balances forward progress, survival, and energy efficiency, with an optional sparse collection term:
$R_{t}=\alpha_{\text{fwd}}\,R^{\text{fwd}}_{t}+\alpha_{\text{energy}}\,R^{\text{energy}}_{t}+\alpha_{\text{alive}}\,R^{\text{alive}}_{t}+K_{t}\,R^{\text{sphere}}_{t},$ (20)
where $\alpha_{\text{fwd}}{=}1$, $\alpha_{\text{energy}}{=}0.005$, $\alpha_{\text{alive}}{=}0.1$, and $K_{t}$ is the number of collectible spheres obtained at step $t$ (if applicable). The forward term encourages task-aligned motion,
$R^{\text{fwd}}_{t}=\begin{cases}\langle v_{t},\ e_{x}\rangle,&\text{flat terrains},\\[2.0pt] \langle v_{t},\ u^{\text{goal}}_{t}\rangle,&\text{mountain terrain},\end{cases}$ (21)
where $v_{t}$ is the base linear velocity, $e_{x}$ is the unit vector along the $x$-axis, and $u^{\text{goal}}_{t}$ is the unit direction from the robot to the goal. The energy term penalizes excessive actuation,
$R^{\text{energy}}_{t}=-\,\|\tau_{t}\|_{2}^{2},$ (22)
where $\tau_{t}$ are realized joint torques. The alive term rewards safe operation,
$R^{\text{alive}}_{t}=1\quad\text{until termination (falls or unrecoverable collisions)}.$ (23)
The per-sphere term $R^{\text{sphere}}_{t}$ is a fixed bonus awarded for each collected item.
4.4 Model Architecture Details
As outlined in Fig. 1, the network uses lightweight encoders, an SSM-based fusion backbone, and compact heads for control. Table 1 summarizes the architectural settings.
Table 1: Architecture settings (compact).
Component Setting
Token width $d$ 128
Proprio encoder 2-layer MLP (256, 256), ReLU
Visual token projection CNN patchify $\rightarrow$ linear to width $d$ (=128)
Mamba backbone $L_{m}{=}2$ stacked SSM layers, residual + LayerNorm
Projection head 2-layer MLP (256, 256), ReLU
4.5 Training Schema
[…]
Table 2: PPO training hyperparameters (shared across methods).
Hyperparameter Value
Episode horizon 1000
Discount factor $\gamma$ 0.99
Nonlinearity ReLU
Samples per iteration 8192
Minibatch size 256
Optimization epochs 3
PPO clip parameter $\epsilon$ 0.2
Policy learning rate $1\times 10^{-4}$
Value learning rate $1\times 10^{-4}$
Optimizer Adam
Policy initialization Standard Gaussian
Domain randomization.
To improve robustness, all methods use the same physics randomization at episode reset, with parameters sampled uniformly from the ranges in Table 3. Unless otherwise noted, the randomized values remain fixed for the entire episode. The overall LocoMamba training process, incorporating PPO, domain randomization, and curriculum learning, is illustrated in Algorithm 1.
Table 3: Domain randomization ranges (sampled at episode reset; held constant per episode).
Parameter Range Units
$K_{P}$ (joint stiffness) [40, 90][40,\ 90] N·m/rad
$K_{D}$ (joint damping) [0.4, 0.8][0.4,\ 0.8] N·m·s/rad
Link inertia $0.5,\ 1.5]\times$ default -
Lateral friction [0.5, 1.25][0.5,\ 1.25] N·s/m
Body mass $0.8,\ 1.2]\times$ default -
Motor friction [0.0, 0.05][0.0,\ 0.05] N·m·s/rad
Motor strength $0.8,\ 1.2]\times$ default -
Sensor latency [0, 0.04][0,\ 0.04] seconds
Visual perturbations.
In addition to physics randomization, we inject lightweight depth noise to simulate saturated returns and minor sensor artifacts. At each step, we sample $K\sim\mathcal{U}\{3,\dots,30\}$ pixel locations per depth frame and set their values to the sensor’s maximum range (salt-like saturation). This perturbation is applied consistently across all methods.
Obstacle-density curriculum.
For obstacle-based scenarios, density starts from an easier setting and linearly ramps to the target distribution over training iterations. Curriculum scheduling is identical across methods.
End-to-end optimization.
Gradients from the PPO objective flow through the projection head, the Mamba fusion backbone, and both encoders. We use advantage normalization and gradient clipping, and update for several epochs per iteration with shuffled minibatches.
Algorithm 1 LocoMamba training with PPO, domain randomization, and curriculum
1:Input: policy $\pi_{\theta}$, value $V_{\phi}$, fusion $f_{\text{fuse}}$, horizon $T$, samples/iter $N_{\text{iter}}$, PPO/GAE hyperparameters
2:for iteration $1,2,\dots$ do
3:  Sample obstacle density from curriculum; set scenario-specific textures/appearance
4:  while collected samples $N_{\text{iter}}$ do
5:   Sample physics params from Table 3; reset environment
6:   for $t=1$ to $T$ do
7:     Observe $o_{t}=\{s^{\text{prop}}_{t},\ \text{depth frames}\}$
8:     Inject depth perturbation: set $K\!\sim\!\mathcal{U}\{3,\dots,30\}$ random pixels $\rightarrow$ max range
9:     Compute fused feature $h_{t}=f_{\text{fuse}}(o_{t})$ $\triangleright$ Encoders + Mamba SSM
10:     Sample action $\tilde{a}_{t}\!\sim\!\mathcal{N}(\mu_{\theta}(h_{t}),\text{diag}(\sigma_{\theta}^{2}(h_{t})))$, set $a_{t}=\tanh(\tilde{a}_{t})$
11:     Step env with $a_{t}$; receive $r_{t}$, $o_{t+1}$, done
12:     Store $(h_{t},a_{t},r_{t},\log\pi_{\theta}(a_{t}|h_{t}),V_{\phi}(h_{t}),\text{done})$
13:     if done then break
14:     end if
15:   end for
16:  end while
17:  Compute advantages with GAE: $\delta_{t}=r_{t}+\gamma V_{\phi}(h_{t+1})-V_{\phi}(h_{t})$, $A_{t}=\sum_{l\geq 0}(\gamma\lambda)^{l}\delta_{t+l}$
18:  Normalize advantages; compute returns $\hat{R}_{t}$
19:  for epoch =1=1 to $E$ do
20:   for minibatch $\mathcal{B}$ do
21:     $\rho_{t}=\frac{\pi_{\theta}(a_{t}|h_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|h_{t})}$ for $t\in\mathcal{B}$
22:     $\mathcal{L}_{\text{clip}}=\mathbb{E}_{\mathcal{B}}\big{[}\min(\rho_{t}A_{t},\ \text{clip}(\rho_{t},1-\epsilon,1+\epsilon)A_{t})\big{]}$
23:     $\mathcal{L}_{V}=\mathbb{E}_{\mathcal{B}}[(V_{\phi}(h_{t})-\hat{R}_{t})^{2}]$,   $\mathcal{H}=\mathbb{E}_{\mathcal{B}}[\mathcal{H}(\pi_{\theta}(\cdot|h_{t}))]$
24:     Update $\theta,\phi$ to minimize $\mathcal{L}_{\text{clip}}+\beta_{V}\mathcal{L}_{V}-\beta_{H}\mathcal{H}$ (with grad clipping)
25:   end for
26:  end for
27:  Advance curriculum schedule
28:end for
5 Experimental Evaluation
5.1 Evaluation Setup
Research Questions
We organize our evaluation around the following research questions (RQs):
  • RQ1 (SOTA performance). Does LocoMamba, which fuses proprioception and depth with Mamba SSM backbone, achieve better locomotion performance than SOTA, as measured by return, success rate, and collision rate?
  • RQ2 (Cross-modal effectiveness). Does encoding both proprioception and depth yield better locomotion performance than using either modality alone?
  • RQ3 (Mamba SSM efficiency). Under a fixed compute budget, does the Mamba-based fusion backbone reduce memory use and latency and reach target performance with fewer updates or less wall-clock time than attention-based or hierarchical fusion?
  • RQ4 (Training robustness). Does the proposed end-to-end PPO training schema improve learning stability and support multi-objective optimization?
Evaluation Metrics
We evaluate policies using (i) mean episode return and two domain-specific metrics: (ii) progress distance, defined as the displacement in meters along the task-aligned target direction, and (iii) collisions per episode, computed over a fixed episode length of 1,000 steps. Collision checks are performed at every control step, and the collision metric is reported only for obstacle-bearing scenarios and for episodes that encounter at least one obstacle.
Baselines
To assess the effectiveness of the proposed LocoMamba, we conduct comparisons with the following methods:
  • Proprio-only. Uses only the 93-D proprioceptive vector; no exteroception.
  • Proprio-Vision-Only. Encodes proprioception and depth with the same encoders as ours; projects depth features to match the proprio feature width and concatenates the two vectors, which are fed directly to the policy/value heads (no sequence fusion).
  • Transformer Proprio-Vision. Self-attention over visual tokens plus a proprio token, representing attention-based cross-modal fusion (Yang et al., 2021).
  • Transformer Vision-only. Following Transformer Proprio-Vision, we also evaluate a transformer-based model using depth-only input.
All agents are trained using the same PPO schema, curriculum, and domain randomization, and are evaluated under identical settings. Where applicable, we reuse the same proprioceptive and depth encoders, as well as the same token width, to minimize confounding factors related to representation size; the only difference lies in the fusion module.
5.2 Performance vs. Baselines on Simulation Scenarios
We evaluate LocoMamba and its variant, Mamba Vision-Only, which uses only vision as input, against established baselines in the Thin Obstacle environment. Figure 4 shows training curves with mean and one standard deviation across seeds. Table 4 reports means and standard deviations over seeds for episode return, collisions per episode, and progress distance.
From Fig. 4, LocoMamba exhibits faster convergence (steeper early learning slope), higher asymptotic return, and lower variance than all baselines. These trends substantiate its SOTA performance, answering RQ1, and indicate stable optimization under the proposed PPO training protocol, addressing RQ4.
Figure 4: Training learning curves on Thin Obstacle. Solid lines denote the mean episode return across seeds; shaded regions indicate $\pm$ one standard deviation.
Table 4: Performance on the Thin Obstacle environment.
Model Architecture Reward Collision Times Distance Moved
Proprio-Only 145.64 $\pm$ 89.55 487.80 $\pm$ 114.13 6.13 $\pm$ 2.49
Transformer Vision-Only 187.45 $\pm$ 93.88 - 6.75 $\pm$ 2.90
Mamba Vision-Only 28.16 $\pm$ 34.47 - 2.92 $\pm$ 1.01
Transformer Proprio-Vision 511.96 $\pm$ 247.30 141.83 $\pm$ 158.47 24.85 $\pm$ 7.34
LocoMamba (Ours) 762.34 $\pm$ 156.53 72.53 $\pm$ 79.47 32.41 $\pm$ 5.23
From Table 4, Mamba Proprio-Vision attains the best overall performance, with higher return ($762.34\pm 156.53$), fewer collisions ($72.53\pm 79.47$), and longer progress distance ($32.41\pm 5.23$) than all baselines. Relative to Transformer Proprio-Vision, it improves return by $48.9\%$, reduces collisions by $48.9\%$, and increases distance by $30.4\%$. Compared with Proprio-Only, it yields a $423.4\%$ gain in return, an $85.1\%$ reduction in collisions, and a $428.7\%$ increase in distance. The two vision-only variants move little on average, so collision counts are not informative and are omitted. These results indicate that cross-modal fusion is essential for obstacle negotiation and that the selective state-space backbone confers a clear advantage under identical training and evaluation protocols. The findings answer RQ1 affirmatively and support the effectiveness of LocoMamba. Moreover, under the proposed end-to-end PPO training protocol with a compact state-centric reward, the agent exhibits stable learning and balanced trade-offs among progress, smoothness, and safety, thereby answering RQ4.
5.3 Ablation Studies
Ablation on modalities.
From Table 4, Transformer Vision-State improves over Proprio-Only by about 251.5% in return, reduces collisions by about 70.9%, and increases progress distance by about 305.4%. It also outperforms Transformer Vision-Only with about 173.1% higher return and about 268.1% greater distance. LocoMamba further widens these margins: relative to Proprio-Only it achieves about 423.4% higher return, about 85.1% fewer collisions, and about 428.7% longer distance; compared with Mamba Vision-Only it delivers roughly 27.1-fold the return and 11.1-fold the distance. Collision counts for vision-only variants are omitted because these policies move too little for the metric to be informative. These results demonstrate the effectiveness of combining proprioception and depth and answer RQ2.
Ablation on the Cross-Modal Fusion Backbone
We compare fusion backbones under matched modality settings. In the proprioception–vision setting, LocoMamba outperforms Transformer Proprio-Vision on all metrics in Table 4. The return rises from $511.96\pm 247.30$ to $762.34\pm 156.53$ (a $48.9\%$ increase), the number of collisions drops from $141.83\pm 158.47$ to $72.53\pm 79.47$ (a $48.9\%$ reduction), and the progress distance increases from $24.85\pm 7.34$ to $32.41\pm 5.23$ (a $30.4\%$ increase). In the vision-only setting, both methods perform poorly relative to cross-modal policies. Mamba Vision-Only remains substantially below Transformer Vision-Only in return and distance, underscoring the importance of proprioception for egocentric perception and control. These comparisons indicate that the Mamba SSM backbone is most beneficial when fusing modalities within a single policy.
To assess efficiency, Table 5 summarizes learning-curve statistics under identical training schedules. In the proprioception–vision setting, the Mamba backbone achieves a higher final reward (737.4 compared with 714.1), a steeper early-learning slope (6.41 per epoch compared with 3.66, a $75\%$ increase), higher learning efficiency (1.44 compared with 1.28), and a larger area under the learning curve per epoch (601.7 compared with 529.8). In the vision-only setting, the Mamba backbone attains a higher final reward (676.0 compared with 577.8) and a steeper early slope (4.22 compared with 3.69), with a comparable area under the curve. Taken together, these results support RQ3: the Mamba-based backbone improves optimization dynamics under the same training budget and confers efficiency benefits that translate into faster learning and stronger final performance when cross-modal fusion is used. In addition, the learning curves in Fig. 4 show faster convergence and lower variance for the Mamba-based models, indicating more stable optimization. This further supports the efficiency advantage of the Mamba SSM backbone (RQ3) and the robustness of the proposed PPO training protocol (RQ4).
Table 5: Learning Efficiency Analysis: Performance and Speed
Model Final Reward Early Learning Slope Learning Efficiency AUC per Epoch
Proprio-Only 27.2 1.69 0.27 36.5
Transformer Vision-Only 577.8 3.69 1.07 448.3
Mamba Vision-Only 676.0 4.22 1.15 443.8
Transformer Proprio-Vision 714.1 3.66 1.28 529.8
LocoMamba (Ours) 737.4 6.41 1.44 601.7
  • Final Reward: average reward over the last 120 epochs ( 2M samples). Early Learning Slope: reward gain per epoch (first 120 epochs). Learning Efficiency: overall reward gain per epoch. AUC: area under the curve per epoch. Best values are in bold.
5.4 Learning Stability Analysis
Figure 5 reports the coefficient of variation (CoV = std/mean) over the last 200 training epochs for the value-function loss and the estimated advantages. Lower is better.
Figure 5: Coefficient of variation over the last 200 epochs. Bars show CoV for the value-function loss (dark) and advantages (light). Lower values indicate more stable optimization.
LocoMamba (Proprio–Vision) attains the lowest CoV on both metrics (value-loss CoV $0.215$, advantage CoV $0.708$), indicating the most stable optimization. Relative to Transformer Proprio–Vision, the value-loss CoV decreases from $0.550$ to $0.215$ (a $\sim\!61\%$ reduction), and the advantage CoV decreases from $0.932$ to $0.708$ (a $\sim\!24\%$ reduction). Vision-only policies exhibit moderate stability but limited displacement, while Proprio-Only shows highly volatile advantages (CoV $2.334$), consistent with partial observability.
These results corroborate the efficiency and stability benefits of the Mamba SSM backbone (RQ3) and demonstrate that the proposed PPO training protocol yields stable learning with balanced updates (RQ4).
5.5 Generalization to Unseen Conditions
We train in the Thin Obstacle environment and evaluate zero-shot on more challenging, previously unseen terrains. Table 6 reports means and standard deviations over seeds. For vision-only policies, collision counts are omitted because the agents move too little for this metric to be informative.
Table 6: Performance on challenging unseen terrains (zero-shot from Thin Obstacle).
Model Architecture Reward Collisions Distance Moved
Proprio-Only 106.45 $\pm$ 74.94 779.40 $\pm$ 119.85 4.94 $\pm$ 1.97
Transformer Vision-Only 151.75 $\pm$ 82.28 - 5.75 $\pm$ 2.68
Mamba Vision-Only 28.53 $\pm$ 29.25 - 2.79 $\pm$ 0.96
Transformer Proprio-Vision 257.38 $\pm$ 431.52 593.73 $\pm$ 135.62 16.45 $\pm$ 8.25
LocoMamba (Ours) 583.76 $\pm$ 154.57 470.27 $\pm$ 108.15 24.04 $\pm$ 6.18
Cross-modal policies generalize substantially better than single-modality baselines. Relative to Proprio-Only, LocoMamba increases return by approximately $448\%$ (from $106.45$ to $583.76$), reduces collisions by $39.7\%$ (from $779.40$ to $470.27$), and raises progress distance by $386.6\%$ (from $4.94$ to $24.04$). Compared with Transformer Proprio-Vision, LocoMamba achieves a $126.8\%$ higher return, $20.8\%$ fewer collisions, and a $46.1\%$ longer distance. The vision-only variants yield low displacement, indicating weak generalization without proprioceptive anchoring. These out-of-distribution results show that combining proprioception with depth is effective for generalization and that the Mamba SSM improves cross-modal fusion under distribution shift.
6 Conclusion
This paper presented LocoMamba, a vision-driven cross-modal reinforcement learning framework for quadrupedal locomotion. The method embeds proprioceptive state with a lightweight MLP and patchifies depth images with a compact CNN, then fuses the resulting tokens using stacked Mamba selective state-space layers. Policies and values are optimized end to end with PPO under terrain and appearance randomization and an obstacle-density curriculum, guided by a compact state-centric reward.
Across challenging simulated environments, LocoMamba achieved higher returns, higher success rates, fewer collisions, and longer progress distances than strong baselines that include proprioception-only controllers, depth-only controllers, recurrent fusion, and Transformer-based fusion. Learning-curve analyses and efficiency metrics showed faster convergence and improved sample efficiency under the same compute budget, consistent with the near-linear scaling of the selective state-space backbone. Stability indicators, including variability of advantages and value loss, further confirmed reliable optimization.
Ablation studies established the importance of combining proprioception with depth and showed that the Mamba backbone is particularly effective for cross-modal fusion. The proposed PPO training protocol, which combines domain randomization, curriculum scheduling, and a compact reward, produced stable learning and balanced trade-offs among progress, smoothness, and safety. Together, these results answer the research questions on overall performance, efficiency, modality design, and training robustness in the affirmative.
Due to current budgetary and hardware-access constraints, we have not conducted real-world experiments. When resources permit, we will deploy LocoMamba on a quadruped platform and assess sim-to-real transfer, latency, and safety under field conditions.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgment